A person got 100000 dollar when he retired from service. He deposited some of that amount of money in bank and rest of his money in the post office and got 5400 dollar in total per year as interest .If the rate of simple interest per annum in the bank and in the post office are 5% and 6% respectively then let us write by calculating the money he had deposited in the bank and post office.
Answer by ikleyn(52757)

.
    B +     P = 10000,     (1)   ("Bank + Post = 100000")
0.05B + 0.06P =  5400.     (2)   (interest + interest = total interest)


Simplify by multiplying eq(2) by 100 (both sides):

 B +  P = 100000           (3)
5B + 6P = 540000           (4)


Multiply eq(3) by 5 (both sides). Keep eq(4) as is. You will get

5B + 5P = 500000           (5)
5B + 6P = 540000           (6)


Now subtract eq(5) from eq(6) (both sides). You will get

6P - 5P = 540000-500000,   or

P = 40000.


Answer.  $40000 was invested to the post office at 6% and the rest $60000 was invested to the bank at 5%.


Check.   0.05*60000 + 0.06*40000 = 5400 dollars.   ! Correct !

Solved.   //  I used the  Elimination method  to solve the system of  2  equations.

-----------------
To see other similar solved problems on investment,  look into the lesson
    - Using systems of equations to solve problems on investment
in this site.